OpenStudy (anonymous):

What is the range of f(x) = -3^x + 4? A. the set of real numbers less than 4 B. the set of real numbers greater than 4 C. the set of real numbers less than 0 D. the set of real numbers greater than 0

OpenStudy (jdoe0001):

what's the graph of \(\bf y= -3^x \quad ?\)

OpenStudy (anonymous):

range will be "the set of real numbers less than 4" \[y \in \mathbb{R} : y < 4\]

OpenStudy (jdoe0001):

\(\bf \textit{parent function}\quad -3^x\\ \quad \\ \begin{array}{rrllll} &\textit{flips it upside down}\\ &\downarrow \\ y=&{\color{red}{ -}}3^x&{\color{red}{ +4}}\\ &&\uparrow \\ &&\textit{vertical shift of 4 units} \end{array}\)

OpenStudy (jdoe0001):

ahemm \(\bf \textit{parent function}\quad -3^x\\ \quad \\ \begin{array}{rrllll} &\textit{flips it upside down}\\ &\downarrow\quad \\ y=&{\color{red}{ -}}3^x&{\color{red}{ +4}}\\ &&\uparrow \\ &&\textit{vertical shift of 4 units} \end{array}\) better... there

OpenStudy (jdoe0001):

http://fooplot.com/#W3sidHlwZSI6MCwiZXEiOiIzXngiLCJjb2xvciI6IiNFODE5MTkifSx7InR5cGUiOjEwMDB9XQ-- <--- add the " - " in front of it, press Tab key then add the " +4 " , press Tab key again notice the "transformations"

OpenStudy (jdoe0001):

and thus the range

OpenStudy (anonymous):

You dont need to go into that detail. If you want to know Domain & Range of a function assume some values of x like {1 , 2 , 3 ....} put them in the equation & by getting answers you will have an idea of the range & domain.